ou take the following sound average sound measurements:

Station dBA Duration (minutes)
1 80 1,434
2 91 2
3 102 58

What is the total exposure percentage for the day?
Answer #1

First we convert the actual duration(Cn) into hours .

i.e Cn at station 1 =23.9 hours

2=0.033 hours

3=0.9666

Then we calculate the allowable duration(Tn) for an 8 hour day ,

Tn=8*2^(90-L)/5 ; where L is the dBA for a station,

T1=8*2^(90-80)/5 =32

In excel , T1=8*2^((90-C3)/5)

(Drag this formula for rest of the cells)

Total exposure percentage =100 * (C1/T1+C2/T2+C3/T3)=138.94 =139 %

Actual Duration Allowable Station dBA Duration Cn(in hours) time(Tn) Cn/Tn 32 0.746875 0.0333333336.96440451|0.004786 1.51571